OpenStudy (akonkel):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p = pounds of jelly beans, which equation in simplest form represents the situation described. 0.31p + 1.41 = 1.56 0.31p + 1.41 = 2.1 –0.31p + 2.06 = 1.56 –0.31p + 2.06 = 2.1

OpenStudy (studygurl14):

0.98p + 0.67c = 1.56 p + c = 2.1 c = 2.1 - p 0.98p + 0.67(2.1 - p) = 1.56 0.98p + 1.407 - 0.67p = 1.56 (round 1.407 to 1.41) 0.31p + 1.41 = 1.56

OpenStudy (studygurl14):

Can I have a medal please? (click best answer)

OpenStudy (akonkel):

sure im new sorry not sure how that worked but yes and thanx for the help

OpenStudy (studygurl14):

You're welcome. and WELCOME TO OPENSTUDY!

OpenStudy (akonkel):

thanx and how do i give u a metal

OpenStudy (studygurl14):

click "best response" to the right of my post

OpenStudy (studygurl14):

it is a blue button with white text

OpenStudy (akonkel):

alright thanx got it i have more math q's LONG AND BORING